OpenStudy (anonymous):

Prove That : ___________ Cos2A+1/Sin(2A+B)+(Sec^2+1)=Cos^2A+2SinACos2A

OpenStudy (anonymous):

use the addition and subtraction theorem here and simplify it. It will be easier that way. I dont give answers :) I help you acheive the answer.

myininaya (myininaya):

\[\frac{\cos(2a)+1}{\sin(2a+b)}+\sec^2(a)+1=\cos^2(a)+2\sin(a)\cos(2a) ?\]

OpenStudy (anonymous):

hmmm,I tried every Law extention ,But Useless... I dont need the answer just gimme me the Key .

OpenStudy (anonymous):

Yes @myininaya ,Ty for clearifying .

OpenStudy (anonymous):

It's important to know if myininaya is right about your question?

OpenStudy (anonymous):

No-data Shes right .

OpenStudy (callisto):

Both sides are not equal?

OpenStudy (anonymous):

@Callisto :we should prove the both of them are equal ., I copied the equation from the Book right.

OpenStudy (cwrw238):

cos2a+ 1 = 2cos^2 a sin(2a + b) = sin2a cos b + cos2a sin 2b these might help

OpenStudy (anonymous):

@cwrw238 :Acually I know all law extention ,Useless .:(

OpenStudy (anonymous):

1 can be taken as Cos0.. Does it help???

myininaya (myininaya):

hey so just to be clear you did mean sec^2(a)+1 right?

myininaya (myininaya):

for that one part right?

OpenStudy (cwrw238):

whats strange here is the disappearance of b on RHS

OpenStudy (anonymous):

@myininaya :Yes sry i forgot the a .

OpenStudy (anonymous):

@cwrw238 : I know . If there a B ..It would be Kind easy .

OpenStudy (callisto):

Any other conditions given in the question?

OpenStudy (anonymous):

@Callisto :hmm,Not sure .. But idk thats gonna help or not ,but this question under the topic of Identites belong to the first quad and the sec. quad .

myininaya (myininaya):

I think people (including myself ) agree that this is not an identity because you do not have B anywhere on the other side of the equation.

OpenStudy (anonymous):

I'am really dont know .. Maybe I should be sure of it again . Ty People for your help